OpenStudy (anonymous):

Find the solution to these system of equations: separate x and y values with a comma. x-6y=-13, x=y=-3

OpenStudy (anonymous):

x=y =-3?

OpenStudy (anonymous):

that's what it said sangya

OpenStudy (sawpnilsrabon):

x-6y=-13 (x=y=-3) when y=-3 x-6*-3=-13 x+18=-13 x=-13-18 x=-31 when x=-3 -3-6y=-13 -6y=-13+3 -6y=-10 6y=10 y=10/6 y=1.66 Ans: (x,y)=(-31,1.66)

jimthompson5910 (jim_thompson5910):

sawpnilsrabon that makes no sense at all. You say y = -3 but then y changes to 1.66. The same problem happens with x as well. My guess is that x=y=-3 should be x+y = -3 or maybe x = y - 3

OpenStudy (anonymous):

thx

OpenStudy (sawpnilsrabon):

you are welcome :)

OpenStudy (sawpnilsrabon):

My guess is that x=y=-3 should be x+y = -3 or maybe x = y - 3 your term have no logic, because x and y are meaning separate value.